OpenStudy (anonymous):

Can you answer this? xx6=18 11-k=2 j-1=9 63/i=7 11-u=7 b-1=6 7xs=35 10-p=8 16/y=8 m+5=14 t/4=7 z/6=4

OpenStudy (anonymous):

no we are not robots

OpenStudy (lilg132):

11 - k = 2 -k = 2 - 11 -k = -9 k = 9

OpenStudy (lilg132):

thats one now post the rest one by one

OpenStudy (lilg132):

remember to click good answer :)